OpenStudy (anonymous):

solve the system of equations by the elimination method: 11x-8y=92 9x+5y=6

OpenStudy (anonymous):

how do you eliminate one vairables to solve for the other

OpenStudy (anonymous):

Just rearrange the terms in 1 of the equations and substitute in the other. 11x-8y=92 9x+5y=6 Put the 1st equation as x=(92+8y)/11 Substitute this x in equation 2 and solve for x. And later for y